# Almost device-independent certification of GME states with minimal measurements

### 28 Feb 2024 v1

@@ Appendix B Appendix B: Self-testing Schmidt states
where $d=\min\{d_{A},d_{B_{i}}\}$ and $\boldsymbol{\alpha}$ is a vector composed of the Schmidt coefficients $\alpha_{i}$, such that $10<\alpha_{i}<$ 1 for all $i$ and $\sum_{i=0}^{d-1}\alpha_{i}^{2}=1$. Next, we are going to introduce a steering operator that provides an inequality maximally violated by states of the form (66).
@@
In our scenario, we consider that all the $N-1$ Bobs perform two arbitrary measurements denoted by $B_{i_{0}}$ and $B_{i_{1}}$. Also, we assume that the local dimensions of the state they share are unknown. At the same time, because Alice’s device is trusted, her measurements are completely characterized and the local dimension of the shared state on Alice’s side is known. Alice’s observables are the generalized Pauli $Z$ and $X$ operators
@@
$S_{d}(\boldsymbol{\alpha})=$ $d\sum_{j=1}^{N-1}\sum_{a,b_{j}=0}^{d-1}c_{a,b_{j}}p(a,b_{j}|0,0)+\gamma(\boldsymbol{\alpha})\left(d\sum_{a,\mathbf{b}=0}^{d-1}c_{a,\mathbf{b}}p(a,\mathbf{b}|1,1,\ldots,1)-\sum_{\begin{subarray}{c}i,a=0\\ i\neq a\end{subarray}}^{d-1}\alpha_{i}\frac{p(a|0)}{\alpha_{a}}\right)$ (80)
@@
$S_{d}(\boldsymbol{\alpha})=$ $d\sum_{j=1}^{N-1}\sum_{a,b_{j}=0}^{d-1}c_{a,b_{j}}p(a,b_{j}|0,0)+\gamma(\boldsymbol{\alpha})\left(d\sum_{a,\mathbf{b}=0}^{d-1}c_{a,\mathbf{b}}p(a,\mathbf{b}|1,1,\ldots,1)-\sum_{i=0}^{d-1}\alpha_{i}\sum_{a=0}^{d-1}\frac{p(a|0)}{\alpha_{a}}\right)-(N-2),$ (82)
@@
$\max_{|\psi_{N}\rangle}\langle\psi_{N}|\mathcal{S}_{d}(\boldsymbol{\alpha},N)|\psi_{N}\rangle\leqslant(d-1)(N-1)+\max_{|\psi_{N}\rangle}\langle\psi_{N}|T_{d}(\boldsymbol{\alpha},N)|\psi_{N}\rangle$ (94)
@@
since $\mathcal{S}_{d}$ is an Hermitian operator and every $\lambda_{i}$ is positive. In this expression we use the fact that Re($z$)$\leqslant|z|$ and $(B^{k}_{j|0})^{\dagger}B^{k}_{j|0}=\mathbbm{1}$ for any $k,j$, to obtain
@@
Consequently, if all the parties attain the quantum bound $(d-1)(N-1)+1$ of $\mathcal{S}_{d}(\boldsymbol{\alpha},N)$, the following conditions hold true
@@
[…]
@@
$(1+\gamma(\boldsymbol{\alpha}))+\frac{\gamma(\boldsymbol{\alpha})}{d}\left(\sum_{k=1}^{d-1}\sum_{\begin{subarray}{c}i,j=0\\ i\neq j\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}\omega^{k(l-j)}\right)$ $=(1+\gamma(\boldsymbol{\alpha}))+\frac{\gamma(\boldsymbol{\alpha})}{d}\left(\sum_{\begin{subarray}{c}i=0\\ i\neq l\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{l}}(d-1)-\sum_{\begin{subarray}{c}i,j=0\\ i\neq j,j\neq l\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}\right)$ (190)
@@
$(1+\gamma(\boldsymbol{\alpha}))+\gamma(\boldsymbol{\alpha})\left(\sum_{\begin{subarray}{c}i=0\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{l}}-\frac{1}{d}\sum_{\begin{subarray}{c}i,j=0\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}\right)$ $=\gamma(\boldsymbol{\alpha})\left(1+\frac{1}{d}\sum_{\begin{subarray}{c}i,j=0\\ i\neq j\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}-\frac{1}{d}\sum_{\begin{subarray}{c}i,j=0\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}+\sum_{\begin{subarray}{c}i=0\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{l}}\right),$ (200)
@@
Appendix C Appendix C: Self-test $N$-qubit generalized $W$-States
@@
[…]

### 13 Mar 2026 v2

@@ Appendix B Self-testing Schmidt states
where $d=\min\{d_{A},d_{B_{i}}\}$ and $\boldsymbol{\alpha}$ is a vector composed of the Schmidt coefficients $\alpha_{i}$, such that $0<\alpha_{i}<1$ for all $i$ and $\sum_{i=0}^{d-1}\alpha_{i}^{2}=1$. Next, we are going to introduce a steering operator that provides an inequality maximally violated by states of the form (99).
@@
In our scenario, we consider that all the N1N-1 Bobs perform two arbitrary measurements denoted by $B_{i_{0}}$ and $B_{i_{1}}$. Also, we assume that the local dimensions of the state they share are unknown. At the same time, because Alice’s device is trusted, her measurements are completely characterized and the local dimension of the shared state on Alice’s side is known. Alice’s observables are the generalized Pauli $Z$ and $X$ operators
@@
$S_{d}(\boldsymbol{\alpha})=\$$d\sum_{j=1}^{N-1}\sum_{a,b_{j}=0}^{d-1}c_{a,b_{j}}p(a,b_{j}|0,0)+\gamma(\boldsymbol{\alpha})\left(d\sum_{a,\mathbf{b}=0}^{d-1}c_{a,\mathbf{b}}p(a,\mathbf{b}|1,1,\ldots,1)-\sum_{\begin{subarray}{c}i,a=0\\ i\neq a\end{subarray}}^{d-1}\alpha_{i}\frac{p(a|0)}{\alpha_{a}}\right)$
@@
$S_{d}(\boldsymbol{\alpha})=\$$d\sum_{j=1}^{N-1}\sum_{a,b_{j}=0}^{d-1}c_{a,b_{j}}p(a,b_{j}|0,0)+\gamma(\boldsymbol{\alpha})\left(d\sum_{a,\mathbf{b}=0}^{d-1}c_{a,\mathbf{b}}p(a,\mathbf{b}|1,1,\ldots,1)-\sum_{i=0}^{d-1}\alpha_{i}\sum_{a=0}^{d-1}\frac{p(a|0)}{\alpha_{a}}\right)-(N-2),$ (106)
@@
[…]
@@
[…]
@@
Consequently, if all the parties attain the quantum bound (d1)(N1)+1(d-1)(N-1)+1 of $\mathcal{S}_{d}(\boldsymbol{\alpha},N)$, the following conditions hold true
@@
Assume that the steering inequality (103) is maximally violated when the $0^{\text{th}}$ party Alice chooses the observables $A_{i}$ defined to be $A_{0}=Z$, $A_{1}=X$ and all the other parties choose their observables as $B_{j,i}$ for $i\in\{0,1\}$ and $j=1,2\ldots,N-1$ acting $\mathcal{H}_{B_{j}}$. Let us say that the state which attains this violation is given by $|\psi_{N}\rangle_{A\mathbf{B}E}\in\mathbbm{C}^{d}\otimes\bigotimes_{i=1}^{N-1}\mathcal{H}_{B_{i}}\otimes\mathcal{H}_{E}$, then the following statement holds true for any $d$: $\mathcal{H}_{B_{i}}=\mathcal{H}_{B^{\prime}_{i}}\otimes\mathcal{H}_{B^{\prime\prime}_{i}}$, where $\mathcal{H}_{B^{\prime}_{j}}\equiv\mathbbm{C}^{d}$, $\mathcal{H}_{B^{\prime\prime}_{j}}$ is some finite-dimensional Hilbert space, and there exists a local unitary transformation on Bob’s side $U_{j}:\mathcal{H}_{B_{j}}\rightarrow\mathcal{H}_{B_{j}}$, such that
@@
$(1+\gamma(\boldsymbol{\alpha}))+\frac{\gamma(\boldsymbol{\alpha})}{d}\left(\sum_{k=1}^{d-1}\sum_{\begin{subarray}{c}i,j=0\\ i\neq j\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}\omega^{k(l-j)}\right)$ $=(1+\gamma(\boldsymbol{\alpha}))+\frac{\gamma(\boldsymbol{\alpha})}{d}\left(\sum_{\begin{subarray}{c}i=0\\ i\neq l\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{l}}(d-1)-\sum_{\begin{subarray}{c}i,j=0\\ i\neq j,j\neq l\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}\right)$
@@
$(1+\gamma(\boldsymbol{\alpha}))+\gamma(\boldsymbol{\alpha})\left(\sum_{\begin{subarray}{c}i=0\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{l}}-\frac{1}{d}\sum_{\begin{subarray}{c}i,j=0\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}\right)$ $=\gamma(\boldsymbol{\alpha})\left(1+\frac{1}{d}\sum_{\begin{subarray}{c}i,j=0\\ i\neq j\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}-\frac{1}{d}\sum_{\begin{subarray}{c}i,j=0\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{j}}+\sum_{\begin{subarray}{c}i=0\end{subarray}}^{d-1}\frac{\alpha_{i}}{\alpha_{l}}\right),$
@@
Appendix C Self-test $N$-qubit generalized $W$-States
@@
Assume that the steering inequality (166) is maximally violated when the $0^{\text{th}}$ party Alice chooses the observables $A_{i}$ defined to be $A_{0}=Z$, $A_{1}=X$ and all the other parties choose their observables as $B_{j,i}$ for $i\in\{0,1\}$ and $j=1,2\ldots,N-1$ acting $\mathcal{H}_{B_{j}}$. Let us say that the state which attains this violation is given by $|\psi_{N}\rangle_{A\mathbf{B}E}\in\mathbbm{C}^{2}\otimes\bigotimes_{j=1}^{N-1}\mathcal{H}_{B_{j}}\otimes\mathcal{H}_{E}$, then the following statement holds true: $\mathcal{H}_{B_{j}}=\mathcal{H}_{B^{\prime}_{j}}\otimes\mathcal{H}_{B^{\prime\prime}_{j}}$, where $\mathcal{H}_{B^{\prime}_{j}}\equiv\mathbbm{C}^{2}$, $\mathcal{H}_{B^{\prime\prime}_{j}}$ is some finite-dimensional Hilbert space, and there exists a local unitary transformation on Bob’s side $U_{j}:\mathcal{H}_{B_{j}}\rightarrow\mathcal{H}_{B_{j}}$, such that